Write the Wiedemann Franz law.


1
Expert's answer
2021-03-21T11:26:06-0400

Wiedemann-Franz's law - the statement that the ratio of the thermal conductivity to the electrical conductivity of a metal is proportional to the temperature:


"\frac{k}{\sigma}=(\frac{\pi^2k_b^2}{3e^2})T=L\cdot T" ,


"k" is the thermal conductivity;

"\sigma" is the electrical conductivity;

"L" is the Lorenz number.